OpenStudy (anonymous):

A right triangle has a side length that measures 4 m and a hypotenuse that measures 8.5 m. What is the measure of the other side of the triangle? 7.5 m 8.1 m 9.4 m 9.8 m

OpenStudy (anonymous):

a^2=c^2-b^2=> a^2=8.5^2-4^2=> a=sqrt(56.25) (take +ve a only) => a=7.5 m

OpenStudy (anonymous):

Let x be the other side By pyth. thm, (8.5)^2 = 4^2 + x^2 x=7.5